# Optimization-Free Fast Optimal Control: Bang-Ride Property, Monotonicity, and Applications to Fast Battery Charging

### v1

@@ Abstract
Single-input fast optimal control problems, which aim to achieve the optimal objective as fast as possible, occur in various real-world applications. In the case of fast battery charging, the associated optimal control problem becomes computationally challenging when detailed battery models are used. A recent heuristic optimization-free algorithm can significantly reduce the computational cost and provide an approximate solution, consistent with many heuristic input profiles in practice. These heuristic solutions have several special properties: They follow a bang-ride pattern that always activates a constraint and applies the maximum feasible input. This work investigates when the above properties arise in the optimal input, and ultimately, when the heuristic input profiles satisfy necessary optimality conditions. By exploiting Pontryagin’s maximum principle (PMP), we show that the optimal control is bang-ride under regularity conditions on constraint switching and local controllability of the system. Moreover, the special type of bang-ride behavior, i.e., applying the maximum feasible input, arises under the monotonicity of the system, objective function, and restricted sensitivity of the constraints. These results provide a theoretical justification for a class of charging heuristics and the fast optimization-free algorithm.
@@ 1 INTRODUCTION
Many real-world control tasks can be formulated as fast optimal control problems with a single input, such as the fast charging of batteries [1] or the rapid thawing of frozen cells for cell therapy [2]. In this work, we investigate several control-theoretical questions that arise in the context of fast battery charging. While battery charging serves as our motivation, the theoretical results in this work apply to single-input optimal control in general.
@@
Charging batteries as quickly as possible without violating safety constraints and degrading battery health is crucial for various applications, such as electric cars and power storage. In practice, many input current profiles for fast charging are heuristic with the following property [3]: They first apply the maximum current, and then the current is adjusted to maintain a series of active constraints. For example, the constant-current-constant-voltage (CC-CV) profile first applies the maximum current until reaching the maximum allowed voltage. Then the current is adjusted to maintain the maximum voltage until the charging task is completed. Different charging profiles differ in the set of active constraints.
@@
Model-based optimal control offers an alternative for designing more advanced, degradation-aware charging profiles [1]. Model-based control methods can optimally balance charging speed, safety, and battery lifetime by incorporating a detailed battery model and operation constraints. The bottleneck is the computational cost for solving the resulting optimal control problems, due to complex electrochemical battery models and nonlinear constraints. For example, with the implementation in [4], solving optimal charging problems using the gradient descent algorithm can take approximately between $6$ and $80$ years using one Intel Core i7 CPU [4].
@@
To address the computational bottleneck, a heuristic and optimization-free algorithm, called the hybrid simulation, is proposed in [1] to provide an approximate solution to the optimal control problem. Inspired by the heuristic charging profiles in practice, the hybrid simulation approach to battery charging [1] first applies the maximum current as the active (input) constraint. This active constraint and the battery model form a system of differential algebraic equations (DAEs) that are simulated. Then, whenever a new constraint becomes active during the simulation, such as an upper bound on the voltage, the active constraint is replaced by the new constraint to continue the simulation. This approach uses a single forward simulation to replace optimization, significantly reducing computational time. For example, it provides approximate solutions between $7$ and $30$ hours for the optimal charging problems in [4], compared to between $6$ and $80$ years for solving optimization. However, the obtained approximate solutions can be suboptimal in some settings [1].
@@
We study several theoretical problems motivated by the heuristic battery charging profiles in practice and the hybrid simulation as discussed above. First, many charging profiles, including the solution of the hybrid simulation, always activate a constraint at every time instant, either the maximum input bound or another constraint. This type of solution has been called bang-ride control [5, 6], a generalization of the classical bang-bang control [7] that activates only the input bound. Second, while a general bang-ride control may switch between minimum and maximum input, the charging profiles and the hybrid simulation often apply the maximum feasible input. Finally, the answers to the above theoretical questions should lead to an understanding of when the practical charging profiles and the hybrid simulation are optimal. We summarize the above questions as:
@@
In the above questions, the optimal control in the second question is a special type of the general bang-ride control in the first question. In the third question, we focus on necessary optimality conditions, instead of sufficient conditions, following PMP.
@@
Related Work: The fast-charging problem has motivated several works to address some of the theoretical questions [5, 6, 8, 9]. Bang-ride control has also been studied in process control [10, 11]. The results in [8, 10, 11] are related only to the first theoretical question, i.e., bang-ride control, generalizing the classical bang-bang theory [7, 12]. However, (pure) state constraints, which do not depend on the input, are not considered in [10], and no proof is provided in [11]. It is well-known in optimal control theory [13, 14] that state constraints require special treatment, compared to mixed constraints that depend on both the input and state.
@@
The result in [5] derives the optimal control analytically for a particular optimal charging problem with a linear constraint and shows that the resulting input is bang-ride. However, Ref. [5] does not provide a general theory for a wide class of control problems. Ref. [6] addresses the second question and observes that the monotonicity of the system, objective, and constraints play important roles. Particularly, if the system is monotonic [15], i.e., a larger input leads to a larger state, and if an objective is minimized and is monotonically decreasing in the state, a maximum input is optimal when there is only the input bound. However, the effect of general constraints is more difficult to capture. The result in [6] shows that the maximum input remains optimal if the nonlinear constraints are increasing in the input and decreasing in the state, i.e., applying larger state and input does not violate the constraints. However, our first question related to general bang-ride control is not considered in [6], and many practical constraints are not decreasing in the state. Finally, Ref. [9] also studies the second question and observes the importance of the monotonicity of the system; however, there are technical errors in the result, as mentioned by its authors [9].
@@
Contribution: We address all three theoretical questions in this work. To obtain bang-ride control, we show that, in addition to the classical condition on local controllability of the system considered also in bang-bang control, the regularity of switchings of state constraints is also important. Particularly, a sufficient condition is to have an input discontinuity when a new state constraint becomes active. Such input discontinuity is indeed observed in the application [5], e.g., where the optimal current jumps when a state constraint is activated. The above result is derived by exploiting the classical regularity conditions in optimal control theory [13]. This result generalizes [10, 12] by including both mixed constraints and state constraints.
For the second question, we first establish a result analogous to the one in [6]; however, we take a different approach via PMP and formulate the problem equivalently as achieving a positive switching function, a function that also appears in classical bang-bang control and determines the switching of inputs. This viewpoint via PMP unifies all our research questions and allows us to address state constraints more formally. More importantly, it enables us to develop further results for linear systems with more general nonlinear constraints that may increase with the state. Here, we show that the maximum feasible input being optimal is equivalent to the external positivity [16, 17], i.e., having a positive output, of a dynamical system, consisting of the costate dynamics from PMP and the switching function as the output. Besides monotonicity of the system, we show that the external positivity can be achieved by further restricting the relative sensitivity of the constraints, i.e., the sensitivity of a constraint to the state relative to its sensitivity to the input.
@@
Combining the above results addresses the third question. We also show the application of the theoretical results to a battery fast-charging problem, formulated for the single-particle battery model [18] and a nonlinear voltage constraint.
@@
Outline: This article is structured as follows. Preliminaries and the problem are introduced in Section 2. Bang-ride control is studied in Section 3. The last two questions are addressed in Section 4 and are extended to more general constraints with linear systems in Section 5. The theory is applied to a battery charging problem in Section 6. All proofs are presented in the Appendix.
@@
Notation: For any matrix $M\in\mathbb{R}^{n\times m}$, $M\geqslant 0$ and $M>0$ denote element-wise inequality, i.e., $M$ only contains non-negative and positive elements, respectively. The symbols $1_{n\times m}$ and $0_{n\times m}$ denote $n\times m$ matrices with only ones and zeros, respectively. The notation $\mathrm{diag}(a_{1},a_{2},\dots,a_{n})$ is a diagonal matrix with diagonal elements $a_{i}$. The optimal trajectory of the optimal control problem is denoted by $(x^{\star},u^{\star})$. For any function $s(x(t),u(t))$, we define shorthand notations $s[t]:=s(x(t),u(t))$ and $s^{\star}[t]:=s(x^{\star}(t),u^{\star}(t))$. For brevity, the partial derivatives $\partial s/\partial x$ and $\partial s/\partial u$ are denoted as $s_{x}$ and $s_{u}$, respectively. A square real matrix is a Metzler matrix if all its off-diagonal elements are non-negative.
@@ 2.1 Nonlinear Systems
The derivation of system-theoretical properties of (1), such as the nonlinear generalizations of controllability/observability, typically involves the operator known as the Lie bracket [19]. Given two functions $f,\bar{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ in class $C^{\infty}$, the Lie bracket is defined as
@@ 2.2 Optimal Control Theory
where the terminal time $t_{f}$ can be fixed or a free optimization variable, which will be specified when necessary. Functions $s:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{c_{s}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{c_{h}}$ denote the mixed (state-input) constraints and the (pure) state constraints, respectively. We assume that $s$ includes an input bound $u\in[u_{\mathrm{min}},u_{\mathrm{max}}]$, and the remaining constraints are called the path constraints. Functions $l,\phi:\mathbb{R}^{n}\to\mathbb{R}$ and $z:\mathbb{R}^{n}\to\mathbb{R}^{c_{z}}$ are the stage cost, terminal cost, and terminal constraint. The cost $l$ is independent of $u$ as we focus on fast optimal control problems that do not penalize the input.
PMP provides the necessary conditions for optimal solutions of (2). In PMP, regularity conditions on the time derivatives of state constraints are typically imposed [14]. Consider the first-order time derivative [^1] When $\dot{h}\equiv 0$, it is necessary to consider higher order time derivatives, see [14, Thm. 6.1]. To focus on the main result of this work, this technical complexity is avoided in this work. of $h$,
@@
where each $\bar{s}_{i}$ depends on $u$. For brevity, $\bar{s}_{i}$ is said to be active, for $i\in\{c_{s}+1,\dots,c_{s}+c_{h}\}$, if and only if (iff) the associated state constraint $h_{i-c_{s}}$ is active.
@@
    Functions $f$, $g$, $\phi$, $s$, $l$, and $z$ are continuously differentiable, and each component of $h$ is in $C^{2}$.
@@
Assumption 2.1b is the standard constraint qualification [14], analogous to constraint qualifications in optimization theory. It also implies that at most one constraint is active at any time, as $u(t)$ is a scalar. A time instant $t$ is called an entry time if some state constraint $h_{i}$ is inactive for a (non-zero) time interval before $t$ and becomes active for a time interval starting at $t$.
@@
where $\lambda_{0}$ is a constant, and $\lambda:[0,t_{f}]\to\mathbb{R}^{n}$, $\mu:[0,t_{f}]\to\mathbb{R}^{c_{s}}$, $\nu:[0,t_{f}]\to\mathbb{R}^{c_{h}}$ are multiplier functions. Moreover, define the state-dependent control set [14]
@@
Then, under Assumption 2.1, PMP [14, Thm. 5.1] states that, given an optimal solution $(x^{\star},u^{\star})$ of (2), there exist $\lambda_{0}\leqslant 0$, functions [^2] The signs of $\lambda_{0}$, $\mu$, $\nu$, $\gamma$, and $\eta$ differ from [14] as we consider minimizing the objective and reversed inequality constraints. $\lambda$ is piecewise absolutely continuous, and $\mu$ and $\nu$ are piecewise continuous. $\lambda$, $\mu$, $\nu$, $\alpha\in\mathbb{R}^{c_{z}}$, $\gamma\in\mathbb{R}^{c_{h}}$, and a vector $\eta(\tau_{i})\in\mathbb{R}^{c_{h}}$ for each time $\tau_{i}$ of discontinuity of $\lambda$ such that, for any $t\in[0,t_{f}]$,
@@
At each entry time [^3] Some literature also considers potential discontinuities at contact time [14], i.e., time $t$ when a state constraint is active but is inactive before and after $t$. As this is not often encountered in the application, we assume no contact time exists for simplicity. $\tau_{i}$, $\lambda_{j}(\tau_{i})$, for some $j\in\{1,\dots,n\}$, may have a discontinuity satisfying
@@ 2.4 Problem Formulation
The system shows that given a fixed initial state, applying a larger input leads to a larger state trajectory. A system with this property is called a monotone system [15] and is also a positive system [24]. Then, increasing the state also decreases the objective $x_{1}(t_{f})$. Therefore, a reasonable guess is that the optimal input applies $u=u_{\mathrm{max}}=1$ and then rides the path constraint (13) if active, i.e., consistent with the solution of the hybrid simulation. This is indeed optimal when $a_{1}=0$ and $a_{2}=-1$, as shown in Fig. 1.
@@
The associated state trajectories in Fig. 3 show that the hybrid simulation initially increases both states; however, the constraint soon becomes active and limits the input, decreasing $x_{1}$ and thus having a suboptimal objective. Instead, the optimal control increases the states later. It avoids early activation of the path constraint and thus leads to a larger $x_{1}$ at the terminal time. Therefore, this performance gap between the hybrid simulation and optimal control is due to the combined effect of system dynamics and the constraint. $\triangle$

### v3

# Bang-Ride Optimal Control: Monotonicity, External Positivity, and Fast Battery Charging

@@ Abstract
This work studies a class of optimal control problems with scalar inputs and general constraints, whose solutions follow a bang-ride pattern that always activates a constraint and enables efficient numerical computation. As a motivating example, fast battery charging leads to computationally demanding optimal control problems when detailed electrochemical models are used. Recently proposed optimization-free heuristics reduce this computational cost while producing input profiles observed in practice, following a bang-ride pattern and applying the maximum feasible input. We investigate when such heuristics satisfy necessary optimality conditions. By leveraging Pontryagin’s maximum principle, we unify and formalize existing insights on the bang-ride structure and on the optimal control attaining the maximum feasible input under monotonicity. We further establish a novel connection between the structured optimal control and the external positivity of the costate dynamics. These results provide a rigorous theoretical foundation for heuristic charging strategies and explain the efficiency of optimization-free algorithms.
@@ 1 INTRODUCTION
Many real-world control tasks can be formulated as optimal control problems with a single input, such as charging batteries [1] or rapid thawing of frozen cells for cell therapy [2]. However, solving the associated optimal control problems is often computationally demanding, requires significant resources and renders real-time implementation on limited platforms infeasible, despite recent advances in numerical optimal control [3]. This work considers optimal control problems whose structure allows one to obtain optimal or near-optimal solutions efficiently, without heavy online optimization. Under appropriate structural properties such as bang-ride behavior or monotonicity, the optimal input can often be determined by simple rules (e.g., applying the maximum feasible input) or validated a posteriori by simulation, enabling solutions that are orders of magnitude faster to compute than with conventional solvers. In particular, we take the fast battery charging problem as a motivating example, as the associated control profiles have these structural properties.
@@
Charging batteries as quickly as possible without violating safety constraints and degrading battery health is crucial for a wide range of applications, such as electric cars and power storage. In practice, many input current profiles for fast charging are heuristic, based on the following concept [4]: They first apply the maximum current, and then the current is adjusted to maintain a series of active constraints. For example, the constant-current-constant-voltage (CC-CV) profile first applies the maximum current until the maximum allowed voltage is reached. The current is then adjusted to maintain the maximum voltage until the charging task is complete. Different charging profiles differ in the set of active constraints.
@@
Model-based optimal control offers an alternative for designing more advanced, degradation-aware charging profiles [1]. Model-based control methods can optimally balance charging speed, safety, and battery lifetime by incorporating a detailed battery model and operation constraints. The bottleneck is the required resources to solve the resulting optimal control problems due to complex electrochemical battery models and nonlinear constraints. For example, with the implementation in [5], solving optimal charging problems using the gradient descent algorithm is estimated to take between $6$ and $80$ years using one Intel Core i7 CPU [5].
@@
To address the computational bottleneck, a heuristic and optimization-free algorithm, called the hybrid simulation, is proposed in [1] to provide an approximate solution to the optimal control problem. Inspired by heuristic charging profiles in practice, the hybrid simulation approach to battery charging [1] first applies the maximum current as the active (input) constraint. This active constraint and the battery model form a system of differential algebraic equations (DAEs) that are simulated. Then, whenever a new constraint becomes active during the simulation, such as an upper bound on the voltage, the active constraint is replaced by the new constraint to continue the simulation. This approach uses a single forward simulation to replace optimization, significantly reducing computational time. For example, it provides approximate solutions orders of magnitude faster than applying optimization solvers [5]. However, the approximate solutions obtained can be suboptimal in some settings [1].
@@
We study several theoretical problems motivated by the heuristic battery charging profiles in practice and hybrid simulation as discussed above. First, many charging profiles, including the solution of the hybrid simulation, activate a constraint at every time instant, either the maximum input bound or another constraint. This type of solution has been called bang-ride control [6, 7], a generalization of the classical bang-bang control [8] that activates only the input bound. Second, while a general bang-ride control may switch between minimum and maximum input, the charging profiles and the hybrid simulation often apply the maximum feasible input. Finally, the answers to the above theoretical questions should lead to an understanding of when the practical charging profiles and the hybrid simulation are optimal. In summary this leads to the following core questions considered in this work:
@@
In the above questions, the optimal control in the second question is a special type of the general bang-ride control in the first question. In the third question, we focus on necessary optimality conditions, instead of sufficient conditions, following Pontryagin’s maximum principle (PMP).
@@
Related Work: The fast-charging problem has motivated several works to address some of the theoretical questions [6, 7, 9, 10]. Bang-ride control has also been studied in process control [11, 12]. The results in [9, 11, 12] are related only to the first theoretical question, i.e., bang-ride control, generalizing the classical bang-bang theory [8, 13]. However, (pure) state constraints, which do not depend on the input, are not considered in [11], and no proof is provided in [12]. It is well known in optimal control theory [14, 15] that state constraints require special treatment, compared to mixed constraints that depend on both the input and the state.
@@
The result in [6] derives the optimal control analytically for a particular optimal charging problem with a linear constraint and shows that the resulting input is bang-ride. However, [6] does not provide a general theory for a wide class of control problems. Ref. [7] addresses the second question and observes that the monotonicity of the system, the objective, and the constraints play important roles. In particular, if the system is monotonic [16], that is, a larger input leads to a larger state, and if an objective is minimized and is monotonically decreasing in the state, a maximum input is optimal when there is only the input bound. However, the effect of general constraints is more difficult to capture. The result in [7] shows that the maximum input remains optimal if the nonlinear constraints are increasing in the input and decreasing in the state, that is, applying larger state and input does not violate the constraints. However, our first question related to general bang-ride control is not considered in [7], and many practical constraints are not decreasing in the state. Finally, [10] also studies bang-ride control and observes the importance of the monotonicity of the system in battery charging problems.
@@
Contribution: We focus on the three theoretical questions described above and demonstrate that PMP provides a unified framework for their analysis. To achieve bang-ride optimal control, while classical results show it is closely connected to nonlinear controllability [11, 12], we incorporate more general constraints and formalize the associated regularity condition on constraint switching, by exploiting classical regularity conditions in optimal control theory [14]. In particular, the regularity condition requires an input discontinuity when a new state constraint becomes active, which is indeed observed in the application [6], e.g., where the optimal current jumps when a state constraint is activated.
For the second question, we establish a result analogous to the ones in [7, 10] but take a different approach via PMP. The result shows that the optimal control has the structure of the maximum feasible input under monotonicity of the system, objective, and constraints. While the theoretical insight is analogous to [7, 10], exploiting PMP allows us to formalize regularity conditions, make an explicit connection to bang-ride control, and make further generalizations.
In particular, the PMP-based approach provides a novel perspective: the structured input is optimal when the dynamic system, composed of the costate dynamics from PMP with the switching function as output, exhibits external positivity [17, 18], i.e., produces positive outputs. When specializing to monotone linear systems, we show that the structured optimal input and external positivity can be achieved by restricting the relative sensitivity of the constraints, i.e., the sensitivity of a constraint to the state relative to its sensitivity to the input.
@@
Combining the above results allows us to address the third question. We show the application of the theoretical results to a fast-charging battery problem, formulated for the single-particle battery model [19] and a nonlinear voltage constraint.
@@
Outline: This article is structured as follows. Preliminaries and the problem are introduced in Section 2. Bang-ride control is studied in Section 3. The last two questions are addressed in Section 4, and external positivity is studied in Section 5. The theory is applied to a battery charging problem in Section 6. All proofs are presented in the Appendix.
@@
Notation: For any matrix $M\in\mathbb{R}^{n\times m}$, $M\geqslant 0$ and $M>0$ denote element-wise inequality, i.e., $M$ only contains non-negative and positive elements, respectively. The symbols $1_{n\times m}$ and $0_{n\times m}$ denote $n\times m$ matrices with only ones and zeros, respectively. The notation $\mathrm{diag}(a_{1},a_{2},\dots,a_{n})$ is a diagonal matrix with diagonal elements $a_{i}$. The optimal trajectory of the optimal control problem is denoted by $(x^{\star},u^{\star})$. For any function $s(x(t),u(t))$, we define the shorthand notation $s[t]:=s(x(t),u(t))$ and $s^{\star}[t]:=s(x^{\star}(t),u^{\star}(t))$. For brevity, the partial derivatives $\partial s/\partial x$ and $\partial s/\partial u$ are denoted as $s_{x}$ and $s_{u}$, respectively. A square real matrix is a Metzler matrix if all its off-diagonal elements are non-negative.
@@ 2.1 Nonlinear Systems
The derivation of system-theoretical properties of (1), such as nonlinear generalizations of controllability/observability, typically involves the operator known as the Lie bracket [20]. Given two functions $f,\bar{f}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ in class $C^{\infty}$, the Lie bracket is defined as
@@ 2.2 Optimal Control Theory
where the terminal time $t_{f}$ can be fixed or a free optimization variable, which will be specified when necessary. The functions $s:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}^{c_{s}}$ and $h:\mathbb{R}^{n}\to\mathbb{R}^{c_{h}}$ denote the mixed (state input) constraints and the (pure) state constraints, respectively. We assume that $s$ includes an input bound $u\in[u_{\mathrm{min}},u_{\mathrm{max}}]$, and the remaining constraints are called the path constraints. The functions $l,\phi:\mathbb{R}^{n}\to\mathbb{R}$ and $z:\mathbb{R}^{n}\to\mathbb{R}^{c_{z}}$ are the stage cost, the terminal cost, and the terminal constraint. Cost $l$ is independent of $u$ in this work.
PMP provides necessary conditions for optimal solutions of (2). In PMP, regularity conditions on the time derivatives of state constraints are typically imposed [15]. Consider the first-order time derivative [^1] When $\dot{h}\equiv 0$, it is necessary to consider higher order time derivatives, see [15, Thm. 6.1]. To focus on the main result of this work, this technical complexity is avoided in this work. of $h$,
@@
where each $\bar{s}_{i}$ depends on $u$. For brevity, $\bar{s}_{i}$ is said to be active for $i\in\{c_{s}+1,\dots,c_{s}+c_{h}\}$ if and only if (iff) the associated state constraint $h_{i-c_{s}}$ is active.
@@
    The functions $f$, $g$, $\phi$, $s$, $l$, and $z$ are continuously differentiable, and each component of $h$ is in $C^{2}$.
@@
Assumption 2.1b is the standard constraint qualification [15], analogous to constraint qualifications in optimization theory. It implies that at most one constraint is active at any time, as $u(t)$ is a scalar. A time instant $t$ is called an entry time if some state constraint $h_{i}$ is inactive for a (non-zero) time interval before $t$ and becomes active for a time interval starting at $t$.
@@
where $\lambda_{0}$ is a constant and $\lambda:[0,t_{f}]\to\mathbb{R}^{n}$, $\mu:[0,t_{f}]\to\mathbb{R}^{c_{s}}$, $\nu:[0,t_{f}]\to\mathbb{R}^{c_{h}}$ are multiplier functions. Moreover, define the state-dependent control set [15]
@@
Then, under Assumption 2.1, PMP [15, Thm. 5.1] states that, given an optimal solution $(x^{\star},u^{\star})$ of (2), there exist $\lambda_{0}\leqslant 0$, functions [^2] The signs of $\lambda_{0}$, $\mu$, $\nu$, $\gamma$, and $\eta$ differ from [15] as we consider minimizing the objective and reversed inequality constraints. $\lambda$ is piecewise absolutely continuous, and $\mu$ and $\nu$ are piecewise continuous. $\lambda$, $\mu$, $\nu$, $\alpha\in\mathbb{R}^{c_{z}}$, $\gamma\in\mathbb{R}^{c_{h}}$, and a vector $\eta(\tau_{i})\in\mathbb{R}^{c_{h}}$ for each time $\tau_{i}$ of discontinuity of $\lambda$ such that, for any $t\in[0,t_{f}]$,
@@
At each entry time [^3] Some literature also considers potential discontinuities at contact time [15], i.e., time $t$ when a state constraint is active but is inactive before and after $t$. As this is not often encountered in the application, we assume that there is no contact time for simplicity. $\tau_{i}$, $\lambda_{j}(\tau_{i})$, for some $j\in\{1,\dots,n\}$, may have a discontinuity satisfying
@@ 2.4 Problem Formulation
The system shows that given a fixed initial state, applying a larger input leads to a larger state trajectory. Such system is called a monotone system [16] and is also a positive system [25]. Then, increasing the state also decreases the objective $x_{1}(t_{f})$. Therefore, a reasonable guess is that the optimal input applies $u=u_{\mathrm{max}}=1$ and then rides the path constraint (13) if active, i.e., consistent with the solution of the hybrid simulation. This is in fact optimal when $a_{1}=0$ and $a_{2}=-1$, as shown in Fig. 1.
@@
The associated state trajectories in Fig. 3 show that the hybrid simulation initially increases both states; however, the constraint soon becomes active and limits the input, decreasing $x_{1}$ and thus having a suboptimal objective. Instead, the optimal control increases the states later, avoiding early activation of the path constraint and leading to a larger $x_{1}$ at the terminal time. Therefore, this performance gap between the hybrid simulation and optimal control is due to the combined effect of system dynamics and the constraint. $\triangle$